# Online Gaussian Process Distance Field with VDB Structure

### v2

@@ Abstract
Robots reason about the environment through dedicated representations. Popular choices for dense representations exploit Truncated Signed Distance Functions (TSDF) and Octree data structures. However, TSDF is a projective signed distance obtained directly from depth measurements that overestimates the Euclidean distance. Octrees, despite being memory efficient, require tree traversal and can lead to increased runtime in large scenarios. Other representations based on Gaussian Process (GP) distance fields are appealing due to their probabilistic and continuous nature, but the computational complexity is a concern. In this paper, we present an online efficient mapping framework that seamlessly couples GP distance fields and the fast-access VDB data structure. This framework incrementally builds the Euclidean distance field and fuses other surface properties, like intensity or colour, into a global scene representation that can cater for large-scale scenarios. The key aspect is a latent Local GP Signed Distance Field (L-GPDF) contained in a local VDB structure that allows fast queries of the Euclidean distance, surface properties and their uncertainties for arbitrary points in the field of view. Probabilistic fusion is then performed by merging the inferred values of these points into a global VDB structure that is efficiently maintained over time. After fusion, the surface mesh is recovered, and a global GP Signed Distance Field (G-GPDF) is generated and made available for downstream applications to query accurate distance and gradients. A comparison with the state-of-the-art frameworks shows superior efficiency and accuracy of the inferred distance field and comparable reconstruction performance. The accompanying code will be publicly available [^3] https://github.com/UTS-RI/VDB_GPDF.
Keywords: Gaussian Process Distance Field, OpenVDB, Euclidean Distance Field, Gradient Field, Mapping.
@@ I Introduction
Recent frameworks based on distance field representations that fulfil some of the above-mentioned requirements have been proposed in the robotics literature [1, 2, 3, 4, 5]. In general, these frameworks are dense and efficient enough to run online, can deal up to a certain point with dynamic objects and are catered to any depth sensor. Most of them incrementally build a projective [1, 3] or non-projective [5] Truncated Signed Distance Field (TSDF) or occupancy map [2, 4] and some can then recover an approximation of the Euclidean Signed Distance Field (ESDF[1, 2, 5, 4]. Despite being dense volumetric representations, these frameworks do not rely on generative models thus the representation is maintained at the resolution that was built and only contains information where the data was observed.
On the other hand, Gaussian Process (GP)-based framework proposed in our previous work [6, 7, 8] are probabilistic generative models that are capable of building accurate Euclidean Distance Field (EDF)s. GP-based frameworks, however, suffer from high computational complexity. To solve this issue and allow online incremental mapping works like [9, 8, 10] use the Octree data structure. The time of accessing the Octree grows with the number of nodes, thus suffering from scalability issues for large datasets. As a fast-access data structure, OpenVDB was used recently proposed in VDBFusion [3]. VDBFusion is a very efficient incremental reconstruction framework based on TSDF, which does not recover the EDF.
@@
  • A generative model that efficiently produces the ESDF via a Global GPDF trained based on the estimated surface which enables downstream applications to query accurate distance and gradients.
We evaluate our proposed VDB-GPDF against the state-of-the-art frameworks showing comparable reconstruction performance and superior efficiency and accuracy of the inferred distance field.
@@ II Related Work
Efficient dense volumetric mapping is appealing for the ability to incrementally build a representation of an unseen environment that can be used for localisation, visualisation, navigation and manipulation in particular when the output space is an ESDF. Inspired by [11], frameworks like Voxblox [1] and Voxfield [12] propose solutions to compute the ESDF as a post-processing step from the projective or non-projective TSDF respectively through wavefront propagation using BFS (Breadth-First Search) algorithm. Using occupancy instead of TSDF, FIESTA [2] also obtains the ESDF from the very efficiently built occupancy map using multiple customised data structures. Employing a less customised data structure, VDB-EDT adopts OpenVDB [13, 14] for occupancy mapping and a distance transform function to represent the EDF hierarchically [4]. Other than occupancy mapping, VDBFusion [3] combines OpenVDB with TSDF fusion to generate volumetric dense maps efficiently. However, VDBFusion does not generate the ESDF as an output. VDBblox [5] adopts the OpenVDB structure to maintain high efficiency. However, VDBblox continues to use TSDF fusion and then propagates it to approximate the ESDF as in [12].
Instead of TSDF or occupancy fusion as all these approaches, we propose to fuse directly on ESDF using an L-GPDF, which allows us to query the Euclidean distance with uncertainty for any point in the field of view. The fusion is then done by merging the inferred EDSF values into the global OpenVDB structure. In Table I, we summarise the most popular online dense volumetric mapping frameworks and add the proposed framework to highlight its difference and novelty.
@@
GP-based distance fields have been proposed as a probabilistic and generative model to represent complex environments [9, 6]. In our prior work [6, 7], we proposed a theoretical sound formulation to estimate the Euclidean distance field by applying the logarithmic transformation to a GP formulation. To improve the accuracy of the distance estimate, we later introduced the reverting GP distance field [8]. Although GP-based frameworks are computationally intensive, Octree structures have enabled fast computations in [9, 8] and more recently in [10] for online incremental mapping. However, Octrees expand the nodes overtime and require tree traversal resulting in a less scalable data structure for large scenarios. Recently, methods like the Gaussian surface model or mixture model [15] have been introduced to model an EDF using fused point clouds or batch mapping, but a substantial gap remains in achieving incremental performance. Via the VDB data structure, our proposed approach facilitates efficiency and scalability while maintaining the advantages of the GPDFs.
@@ III-A Gaussian Process Distance Field
The distance field ${\hat{d}}(\mathbf{x}_{*})$ given any location $\mathbf{x}_{*}$ is obtained by applying the reverting function $r$ to the latent field as
${\hat{d}}(\mathbf{x}_{*})=r\left(\hat{o}\left(\mathbf{x}_{*}\right)\right).$ (3)
Considering the square exponential kernel ${{k}\left(\mathbf{x},\mathbf{x}^{\prime}\right)}=\sigma^{2}\exp\left(-\frac{\lVert\mathbf{x}-\mathbf{x}^{\prime}\rVert{}^{2}}{2l^{2}}\right)$, substituting the reverting function of square exponential kernel into Eq. 3 gives us the distance inference at $\mathbf{x}_{*}$,
${\hat{d}}(\mathbf{x}_{*})=\sqrt{-2l^{2}\log\left(\frac{\hat{o}\left(\mathbf{x}_{*}\right)}{\sigma^{2}}\right)}.$ (4)
and variance as
$\hat{v}(\mathbf{x}_{*})=\Lambda\hat{u}(\mathbf{x}_{*})\Lambda^{\top}.$ (5)
where $\Lambda$ is equal to ${l\sigma_{o}^{2}}/({\hat{o}\left(\mathbf{x}_{*}\right)\sqrt{2\log\left({\hat{o}\left(\mathbf{x}_{*}\right)}/{\sigma_{o}^{2}}\right)}})$. As pointed out in [8], the variance of the latent field is propagated through the reverting function using its Jacobian with the linearisation point at the surface. Faraway from the surface, the uncertainty of the estimated distance field grows rapidly. Therefore, the variance is only informative around the surface.
Furthermore, to estimate the gradient of the distance field, we simply apply the differentiation (linear) operator as
@@ III-B OpenVDB Data Structure
The commonly used Octree data structure [9] usually divides the full scene into spatial clusters in a multi-layer tree and treats each cluster separately. However, as the scene grows, the Octree expands with more nodes, and tree traversing gets slower as the access complexity $\mathcal{O}(\log n)$ is related to the number of nodes $n$. OpenVDB data structure (Volumetric, Dynamic grid that shares several characteristics with B+trees) [14, 13] is used in this work to enable efficient, large-scale, dense incremental mapping. This hierarchical tree-like structure allows us to efficiently divide GP into clusters with constant computational complexity to access them.
@@
The memory and time efficiency properties of VDB [13] are highly suitable for our online mapping frameworks. Note that each leaf node in VDB is a proper-sized GP cluster, and we use the leaf iterator to traverse available voxels in each leaf node and then use them to train each GP cluster separately.
@@ IV VDB-GPDF Framework
Depth sensor data is captured as a raw point cloud ${}^{{C}_{{i}}}\mathcal{P}_{\{i\}}$ in sensor frame ${C}_{{i}}$ at current time $i$. The estimated pose is used to project the point cloud from current sensor frame ${{C}_{{i}}}$ to world frame ${W}$ given the homogeneous transformation matrix $\mathbf{T}_{{C}_{{i}}}$ yielding a point cloud $\mathcal{P}_{\{i\}}$ in the world reference frame. The raw points $\mathcal{P}_{\{i\}}$ are voxelised into the voxels denoted as $\mathcal{P}_{v\{i\}}$ via a local hierarchical VDB. Thus, $\mathcal{P}_{v\{i\}}$ is divided into local clusters maintained by the VDB leaf nodes. For each leaf node, the centres of voxels are used as training points to model GPs separately for distance and surface properties. All GPs together form the temporary Local Gaussian Process Signed Distance Field (L-GPDF) along with surface properties. In addition, $\mathcal{P}_{v\{i\}}$ is passed to the testing points generation block. A ray-cast operation from the current sensor origin to each voxel in $\mathcal{P}_{v\{i\}}$ is performed to generate a set of voxels. These voxels are then used as the testing points $\mathcal{P}_{t\{i\}}$ to query the L-GPDF. Let us denote each point in $\mathcal{P}_{t\{i\}}$ as $\mathbf{x_{*}}$. Note that the distance sign of each $\mathbf{x_{*}}$ is computed given the sensor origin and $\mathcal{P}_{v\{i\}}$. For each querying point $\mathbf{x_{*}}$, the closest node in L-GPDF is found to compute the GP distance inference ${\hat{d}}_{t}(\mathbf{x_{*}})$, surface properties $\hat{c}_{t}(\mathbf{x_{*}})$ and uncertainty.
The $\hat{d}_{t}$ and $\hat{c}_{t}$ estimates (mean and uncertainty) at each location in $\mathcal{P}_{t\{i\}}$ are then fused with the existing map maintained by a global VDB grid, which has all fused voxels $\mathcal{P}_{\{0,...,i\}}$. The leaf nodes in the global VDB grid that participated in the fusion procedure are marked as active. Marching cubes algorithm [23] is then performed for the voxels in the active nodes to recover the dense reconstruction. The zero crossing points $\mathcal{P}_{\mathcal{S}\{0,...,i\}}$ in active nodes and previous active nodes are used to generate the global Gaussian Process Signed Distance Field (G-GPDF). Similar to L-GPDF, each leaf node is trained separately. The G-GPDF is available to be queried for any given point in the set of all points the space $\mathcal{P}_{o}$ of downstream applications that require accurate distance $\hat{d}_{o}$ and gradients $\nabla\hat{d}_{o}$, e.g., navigation, path planning, localisation and manipulation.
@@ IV-A Local Gaussian Process Signed Distance Field
We iterate each leaf node of the local VDB for voxels, which are used as training points to formulate each node as a GP separately. All GPs in local VDB consist of the temporary Local Gaussian Process Signed Distance Field (L-GPDF) along with surface properties. As explained in Section. III-A, given locations of voxels, we compute the distance field inference and variance via Eq. 4 and Eq. 5. In addition, given the voxels’ surface properties of each leaf node, we propose to model the surface properties inference:
@@ IV-B Testing Points Generation
Testing points generation method takes $\mathcal{P}_{v\{i\}}$ and sensor origin as input. Here, we adopt a similar idea as the so-called free space carving method [1, 3] to update the distance information for the voxels within the current sensor field of view. Free-space carving has the ability to update the dynamically changing objects in the scene. We propose three main improvements to enhance efficiency while achieving comparable reconstruction results to the conventional free-space carving used in most volumetric mapping methods.
@@
We use $\mathcal{P}_{t\{i\}}$ to query the distance through the L-GPDF. For each $\mathbf{x_{*}}$ in $\mathcal{P}_{t\{i\}}$, we find the closest node in L-GPDF and use the node GP to compute the GP distance inference $\hat{d}_{t}(\mathbf{x_{*}})$, surface property inference $\hat{c}_{t}(\mathbf{x_{*}})$, and variances $\hat{v}_{t}(\mathbf{x_{*}})$ and $\hat{w}_{t}(\mathbf{x_{*}})$. Thanks to the available Euclidean distance inference of L-GPDF, we only query each voxel once to perform the fusion later. This leads to a comparable result of the final reconstruction but more efficiently than TSDF fusion. Note that we also compute the sign of distance given the sensor origin and $\mathcal{P}_{v\{i\}}$.
@@ IV-C Data Fusion
After querying the testing points from the L-GPDF, we have the signed distance inference $\hat{d}_{t}(\mathbf{x_{*}})$, $\hat{c}_{t}(\mathbf{x_{*}})$ and variance $\hat{v}_{t}(\mathbf{x_{*}})$ $\hat{w}_{t}(\mathbf{x_{*}})$ for each $\mathbf{x_{*}}$ in $\mathcal{P}_{t\{i\}}$. Unlike the commonly used TSDF fusion that uses constant or drop-off weights to fuse the projective (or non-projective) distance, we directly query our probabilistic L-GPDF to fuse the distance and surface properties through its mean and variance in the global VDB. Note that the fusion is mostly done around the surface, as our variance is only accurate near the surface, as discussed in Section III.
The global VDB efficiently accesses the previous distance mean and weight at $\mathbf{x_{*}}\in\mathcal{P}_{t\{i\}}$, which are denoted as $\boldsymbol{D}_{i-1}(\mathbf{x_{*}})$ and $\boldsymbol{V}_{i-1}(\mathbf{x_{*}})$ respectively. Then $\boldsymbol{D}_{i}(\mathbf{x_{*}})$ and $\boldsymbol{V}_{i}(\mathbf{x_{*}})$ is estimated given the new queried $\hat{d}_{t}(\mathbf{x_{*}})$ $\hat{v}_{t}(\mathbf{x_{*}})$ of L-GPDF. We then perform the distance fusion following the standard weighted sum as in [24],
$\boldsymbol{D}_{i}(\mathbf{x_{*}})$ $=\frac{\boldsymbol{V}_{i-1}(\mathbf{x_{*}})\cdot\boldsymbol{D}_{i-1}(\mathbf{x_{*}})+(1-\hat{v}_{t}(\mathbf{x_{*}}))\cdot\hat{d}_{t}(\mathbf{x_{*}})}{\boldsymbol{V}_{i-1}(\mathbf{x_{*}})+(1-\hat{v}_{t}(\mathbf{x_{*}}))}$ (8)
$\boldsymbol{V}_{i}(\mathbf{x_{*}})$ $=\boldsymbol{V}_{i-1}(\mathbf{x_{*}})+(1-\hat{v}_{t}(\mathbf{x_{*}})).$
Similarly, the surface properties are fused using
$\boldsymbol{C}_{i}(\mathbf{x_{*}})$ $=\frac{\boldsymbol{W}_{i-1}(\mathbf{x_{*}})\cdot\boldsymbol{C}_{i-1}(\mathbf{x_{*}})+(1-\hat{w}_{t}(\mathbf{x_{*}}))\cdot\hat{c}_{t}(\mathbf{x_{*}})}{\boldsymbol{W}_{i-1}(\mathbf{x_{*}})+(1-\hat{w}_{t}(\mathbf{x_{*}}))}$ (9)
$\boldsymbol{W}_{i}(\mathbf{x_{*}})$ $=\boldsymbol{W}_{i-1}(\mathbf{x_{*}})+(1-\hat{w}_{t}(\mathbf{x_{*}})).$

### v3

@@ Abstract
Robots reason about the environment through dedicated representations. Popular choices for dense representations exploit Truncated Signed Distance Functions (TSDF) and Octree data structures. However, TSDF provides a projective or non-projective signed distance obtained directly from depth measurements that overestimate the Euclidean distance. Octrees, despite being memory efficient, require tree traversal and can lead to increased runtime in large scenarios. Other representations based on the Gaussian Process (GP) distance fields are appealing due to their probabilistic and continuous nature, but the computational complexity is a concern. In this paper, we present an online efficient mapping framework that seamlessly couples GP distance fields and the fast-access OpenVDB data structure. The key aspect is a latent Local GP Signed Distance Field (L-GPDF) contained in a local VDB structure that allows fast queries of the Euclidean distance, surface properties and their uncertainties for arbitrary points in the field of view. Probabilistic fusion is then performed by merging the inferred values of these points into a global VDB structure that is efficiently maintained over time. After fusion, the surface mesh is recovered, and a global GP Signed Distance Field (G-GPDF) is generated and made available for downstream applications to query accurate distance and gradients. A comparison with the state-of-the-art frameworks shows superior efficiency and accuracy of the inferred distance field and comparable reconstruction performance. https://github.com/UTS-RI/VDB_GPDF
Index Terms:
Gaussian Process Distance Field, OpenVDB, Euclidean Distance Field, Gradient Field, Mapping.
@@ I Introduction
Recent frameworks based on distance field representations that fulfil some of the above-mentioned requirements have been proposed in the robotics literature [1, 2, 3, 4, 5]. In general, these frameworks are dense and efficient enough to run online, can deal up to a certain point with dynamic objects and are catered to any depth sensor. Most of them incrementally build a projective [1, 3] or non-projective [5] Truncated Signed Distance Field (TSDF) or occupancy map [2, 4] and some can then recover an approximation of the Euclidean Signed Distance Field (ESDF) [1, 2, 5, 4]. Despite being dense volumetric representations, these frameworks do not rely on generative models thus the representation is maintained at the resolution that was built and only contains information where the data was observed.
On the other hand, Gaussian Process (GP) framework proposed in our previous work [6, 7, 8] are probabilistic generative models that are capable of building accurate Euclidean Distance Field (EDF)s. GP-based frameworks, however, suffer from high computational complexity. To solve this issue and allow online incremental mapping works like [9, 8, 10] use the Octree data structure. The time of accessing the Octree grows with the number of nodes, thus suffering from scalability issues for large datasets. As a fast-access data structure, OpenVDB [11] was used recently proposed in VDBFusion [3], an efficient incremental reconstruction framework based on TSDF, however, does not produce the EDF.
@@
  • A generative model that efficiently produces the ESDF via a global GPDF trained based on the estimated surface which enables downstream applications to query accurate distance and gradients. Our code is publicly available [^1] https://github.com/UTS-RI/VDB_GPDF.
We evaluate our VDB-GPDF against the state-of-the-art frameworks showing comparable reconstruction performance and superior efficiency and accuracy of the distance field especially in large-scale datasets.
@@ II Related Work
Efficient dense volumetric mapping is appealing for the ability to incrementally build a representation of an unseen environment that can be used for localisation, visualisation, navigation and manipulation in particular when the output space is an ESDF. Frameworks like Voxblox [1] and Voxfield [12] propose solutions to compute the ESDF as a post-processing step from the projective or non-projective TSDF respectively through wavefront propagation using BFS (Breadth-First Search) algorithm. Using occupancy instead of TSDF, FIESTA [2] also obtains the ESDF from the occupancy map using multiple customised data structures. Employing a less customised data structure, VDB-EDT adopts OpenVDB [13, 11] for occupancy mapping and a distance transform function to represent the EDF hierarchically [4]. Other than occupancy mapping, VDBFusion [3] combines OpenVDB with TSDF fusion to generate volumetric dense maps efficiently. However, VDBFusion does not generate the ESDF as an output. VDBblox [5] adopts the OpenVDB structure to maintain efficiency. However, VDBblox continues to use TSDF fusion and then propagates it to approximate the ESDF as in [12].
Instead of TSDF or occupancy fusion as all these approaches, we propose to fuse directly on ESDF using an L-GPDF, which allows us to query the Euclidean distance with uncertainty for any point in the field of view. The fusion is then done by merging the inferred EDSF values into the global OpenVDB structure. In Table. I, we summarise the most popular online dense volumetric mapping frameworks and add the proposed framework to highlight its difference and novelty.
@@
GP-based distance fields have been proposed as a probabilistic and generative model to represent complex environments [9, 6]. In our prior work [6, 7], we proposed a theoretical sound formulation to estimate the Euclidean distance field by applying the logarithmic transformation to a GP formulation. To improve the accuracy of the distance estimate, we later introduced the reverting GP distance field [8]. Although GP-based frameworks are computationally intensive, Octree structures have enabled fast computations in [9, 8] and more recently in [10] for online incremental mapping. However, Octrees expand the nodes over time and require tree traversal, resulting in a less scalable data structure for large scenarios. Recently, methods like the Gaussian mixture model [14] have been introduced to model an EDF using fused point clouds or batch mapping, but a substantial gap remains in achieving incremental performance. Via the VDB data structure, our proposed approach facilitates efficiency and scalability while maintaining the advantages of the GPDFs.
@@ III-A Gaussian Process Distance Field
The distance field ${\hat{d}}(\mathbf{x}_{*})$ given any location $\mathbf{x}_{*}$ is obtained by applying the reverting function $r$ to the latent field as ${\hat{d}}(\mathbf{x}_{*})=r\left(\hat{o}\left(\mathbf{x}_{*}\right)\right)$. Considering the unscaled square exponential kernel ${{k}\left(\mathbf{x},\mathbf{x}^{\prime}\right)}=\exp\left(-\frac{\lVert\mathbf{x}-\mathbf{x}^{\prime}\rVert{}^{2}}{2l^{2}}\right)$, with the kernel’s length scale $l$, the reverting function is defined as
${\hat{d}}(\mathbf{x}_{*})=\sqrt{-2l^{2}\log\left(\hat{o}\left(\mathbf{x}_{*}\right)\right)}.$ (3)
As explained in [8], the resulting distance field is not a GP due to the non-linear reverting operation. Thus its variance is not readily available throughout $\mathbb{R}^{3}$. By simply propagating the latent field’s uncertainty through the first-order Taylor expansion of the reverting function as
$\hat{v}(\mathbf{x}_{*})=\Lambda\hat{u}(\mathbf{x}_{*})\Lambda^{\top},\text{with}\ \Lambda=\frac{-l}{\hat{o}(\mathbf{x}_{*})\sqrt{-2\log\left(\hat{o}(\mathbf{x}_{*})\right)}},$ (4)
the distance uncertainty away from the surface grows indefinitely. However, close to the surface, as required here for the fusion process, $\hat{v}(\mathbf{x}_{*})$ provides a measurement of the uncertainty [^2] As stated in [8], the latent field is not bounded between 0 and 1. The value of $\hat{o}(\mathbf{x}_{*})$ is capped to prevent any singularity on ${\hat{d}}(\mathbf{x}_{*})$ and $\hat{v}(\mathbf{x}_{*})$..
To estimate the gradient of the distance field, we simply apply the differentiation (linear) operator as
@@ III-B OpenVDB Data Structure
OpenVDB (Volumetric, Dynamic grid that shares several characteristics with B+trees) is a hybrid data structure. It splits the space into hash tables and stores a hierarchical tree inside each table, resulting in sharing the advantages of the hashing method. OpenVDB [11, 13] is used to enable efficient, large-scale, dense incremental mapping. This hierarchical tree-like structure allows us to efficiently divide GP into clusters with constant computational complexity to access them.
@@
The memory and time efficiency properties of VDB [13] are highly suitable for our online mapping frameworks. We treat each leaf node in VDB as a GP cluster, and a leaf iterator is used to traverse voxels and train GP clusters separately.
@@ IV VDB-GPDF Framework
Depth sensor data is captured as a raw point cloud ${}^{{C}_{{i}}}\mathcal{P}_{\{i\}}$ in sensor frame ${C}_{{i}}$ at current time $i$. The estimated pose is used to project the point cloud from current sensor frame ${{C}_{{i}}}$ to world frame ${W}$ given the homogeneous transformation matrix $\mathbf{T}_{{C}_{{i}}}$ yielding a point cloud $\mathcal{P}_{\{i\}}$ in the world reference frame. The raw points $\mathcal{P}_{\{i\}}$ are voxelised into the voxels denoted as $\mathcal{P}_{v\{i\}}$ via a local hierarchical VDB. Thus, $\mathcal{P}_{v\{i\}}$ is divided into local clusters maintained by the VDB leaf nodes. For each leaf node, the centres of voxels are used as training points to model GPs separately for distance and surface properties. All GPs together form the temporary L-GPDF along with surface properties. In addition, $\mathcal{P}_{v\{i\}}$ is passed to the testing points generation block. A ray-cast operation from the current sensor origin to each voxel in $\mathcal{P}_{v\{i\}}$ is performed to generate a set of voxels. These voxels are then used as the testing points $\mathcal{P}_{t\{i\}}$ to query the L-GPDF. Let us denote each point in $\mathcal{P}_{t\{i\}}$ as $\mathbf{x_{*}}$. Note that the distance sign of each $\mathbf{x_{*}}$ is computed given the sensor origin and $\mathcal{P}_{v\{i\}}$. For each querying point $\mathbf{x_{*}}$, the closest node in L-GPDF is found to compute the GP distance inference ${\hat{d}}_{t}(\mathbf{x_{*}})$, surface properties $\hat{c}_{t}(\mathbf{x_{*}})$ and uncertainty.
The $\hat{d}_{t}$ and $\hat{c}_{t}$ estimates (mean and uncertainty) at each location in $\mathcal{P}_{t\{i\}}$ are then fused with the existing map maintained by a global VDB grid, which has all fused voxels $\mathcal{P}_{\{0,...,i\}}$. The leaf nodes in the global VDB grid that participated in the fusion procedure are marked as active. Marching cubes algorithm [22] is then performed for the voxels in the active nodes to recover the dense reconstruction. The zero crossing points $\mathcal{P}_{\mathcal{S}\{0,...,i\}}$ in active nodes and previous active nodes are used to generate the G-GPDF. Similar to L-GPDF, each leaf node is trained separately. The G-GPDF is available to be queried for any given points $\mathcal{P}_{o}$ for downstream applications that require accurate distance $\hat{d}_{o}$ and gradients $\nabla\hat{d}_{o}$, e.g., navigation, path planning, localisation and manipulation.
@@ IV-A Local Gaussian Process Signed Distance Field
We iterate each leaf node of the local VDB for voxels, which are used as training points to formulate each node as a GP separately. All GPs in local VDB consist of the temporary L-GPDF along with surface properties. As explained in Section. III-A, given locations of voxels, we compute the distance field inference and variance via Eq. 3 and Eq. 4. In addition, given the voxels’ surface properties of each leaf node, we propose to model the surface properties inference:
@@ IV-B Testing Points Generation
Testing points generation method takes $\mathcal{P}_{v\{i\}}$ and sensor origin as input. Here, we adopt a similar idea as the so-called free space carving method [1, 3] to update the distance information for the voxels within the current sensor field of view. Free-space carving has the ability to update the dynamically changing objects in the scene. We update voxels with previously and newly observed surfaces. We propose three main improvements to enhance efficiency while achieving comparable reconstruction results to the conventional free-space carving used in most volumetric mapping methods.
@@
We use $\mathcal{P}_{t\{i\}}$ to query the distance through the L-GPDF. For each $\mathbf{x_{*}}$ in $\mathcal{P}_{t\{i\}}$, we find the closest node in L-GPDF and use the node GP to compute the GP distance inference $\hat{d}_{t}(\mathbf{x_{*}})$, surface property inference $\hat{c}_{t}(\mathbf{x_{*}})$, and variances $\hat{v}_{t}(\mathbf{x_{*}})$ and $\hat{w}_{t}(\mathbf{x_{*}})$. Thanks to the available Euclidean distance inference of L-GPDF, we only query each voxel once to perform the fusion later. This leads to a comparable result of the final reconstruction but more efficiently than TSDF fusion. Note that we also compute the distance sign given the sensor origin.
@@ IV-C Data Fusion
After querying the testing points from the L-GPDF, we have the signed distance inference $\hat{d}_{t}(\mathbf{x_{*}})$, $\hat{c}_{t}(\mathbf{x_{*}})$ and variance $\hat{v}_{t}(\mathbf{x_{*}})$ $\hat{w}_{t}(\mathbf{x_{*}})$ for each $\mathbf{x_{*}}$ in $\mathcal{P}_{t\{i\}}$. Unlike the commonly used TSDF fusion that uses constant or drop-off weights to fuse the projective (or non-projective) distance, we directly query our probabilistic L-GPDF to fuse the distance and surface properties through its mean and variance in the global VDB. Note that the fusion is mostly done around the surface, as discussed in Section III.
The global VDB efficiently accesses the previous distance mean and variance at $\mathbf{x_{*}}\in\mathcal{P}_{t\{i\}}$, which are denoted as $\boldsymbol{D}_{i-1}(\mathbf{x_{*}})$ and $\boldsymbol{V}_{i-1}(\mathbf{x_{*}})$ respectively. Then new $\boldsymbol{D}_{i}(\mathbf{x_{*}})$ and $\boldsymbol{V}_{i}(\mathbf{x_{*}})$ are estimated given the queried $\hat{d}_{t}(\mathbf{x_{*}})$ $\hat{v}_{t}(\mathbf{x_{*}})$ of L-GPDF. We then perform fusion by using the standard weighting average,
$\boldsymbol{D}_{i}(\mathbf{x_{*}})=\alpha\cdot\hat{d}_{t}(\mathbf{x_{*}})+\beta\cdot\boldsymbol{D}_{i-1}(\mathbf{x_{*}}),\ \boldsymbol{V}_{i}(\mathbf{x_{*}})=\gamma.$ (7)
where the probabilistic weights are given by,
$\alpha$ $=\frac{\boldsymbol{V}_{i-1}(\mathbf{x_{*}})}{\boldsymbol{V}_{i-1}(\mathbf{x_{*}})+\hat{v}_{t}(\mathbf{x_{*}})},\ \beta=\frac{\hat{v}_{t}(\mathbf{x_{*}})}{\boldsymbol{V}_{i-1}(\mathbf{x_{*}})+\hat{v}_{t}(\mathbf{x_{*}})},$ (8)
$\gamma$ $=(\boldsymbol{V}_{i-1}(\mathbf{x_{*}})^{-1}+\hat{v}_{t}(\mathbf{x_{*}})^{-1})^{-1}.$
Alternatively, we can also use our variance to update the weights following [23],
$\alpha$ $=\frac{1-\hat{v}_{t}(\mathbf{x_{*}})}{\boldsymbol{V}_{i-1}(\mathbf{x_{*}})+1-\hat{v}_{t}(\mathbf{x_{*}})},\ \beta=\frac{\boldsymbol{V}_{i-1}(\mathbf{x_{*}})}{\boldsymbol{V}_{i-1}(\mathbf{x_{*}})+1-\hat{v}_{t}(\mathbf{x_{*}})},$ (9)
$\gamma$ $=\boldsymbol{V}_{i-1}(\mathbf{x_{*}})+1-\hat{v}_{t}(\mathbf{x_{*}}).$
Similarly, the surface properties are fused using the same expressions as above but with their respective variance $\hat{w}_{t}(\mathbf{x_{*}})$.